1. A population consists of six values (6, 9, 12, 15, 18, and 21).


a. Select a random sample of size 3. Explain the random sampling that you used.


b. How many possible samples can be drawn?


c. List all possible samples and compute the mean of each sample.


d. Construct a frequency distribution of the sample means obtained in step 2 including 𝑥̅; 𝑓; 𝑃(𝑥̅); ̅𝑥 ⋅ 𝑃(𝑥̅); ̅𝑥 2 ⋅ 𝑃(𝑥̅); Σ𝑃(𝑥̅); Σ𝑥̅⋅ 𝑃(𝑥̅) 𝑎𝑛𝑑 Σ𝑥̅ 2 ⋅ 𝑃(𝑥̅).


1
Expert's answer
2022-05-12T09:21:42-0400

We have population values 6,9,12,15,18,21, population size N=6 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{6+9+12+15+18+21}{6}=13.5"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{1}{6}(56.25+20.25+2.25"


"+2.25+20.25+6.25)=26.25"

"\sigma=\sqrt{\sigma^2}=\sqrt{26.25}\approx5.123475"

a. Select a random sample of size 3 without replacement. We have a sample distribution of sample mean.


b. The number of possible samples which can be drawn without replacement is "^{N}C_n=^{6}C_3=20."


c.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 6,9,12 & 9 \\
 \hdashline
 2 & 6,9,15 & 10 \\
 \hdashline
 3 & 6,9,18 & 11 \\
 \hdashline
 4 & 6,9,21 & 12 \\
 \hdashline
 5 & 6,12,15 & 11 \\
 \hdashline
 6 & 6,12,18 & 12 \\
 \hdashline
 7 & 6,12,21 & 13 \\
 \hdashline
 8 & 6,15,18 & 13 \\
 \hdashline
 9 & 6,15,21 & 14 \\
 \hdashline
 10 & 6,18,21 & 15 \\
 \hdashline 
 11 & 9,12,15 & 12 \\
 \hdashline 
 12 & 9,12,18 & 13 \\
 \hdashline 
 13 & 9,12,21 & 14 \\
 \hdashline 
 14 & 9,15,18 & 14 \\
 \hdashline 
 15 & 9,15,21 & 15 \\
 \hdashline 
 16 & 9,18,21 & 16 \\
 \hdashline 
 17 & 12,15,18 & 15 \\
 \hdashline 
 18 & 12,15,21 & 16 \\
 \hdashline 
 19 & 12,18,21 & 17 \\
 \hdashline 
 20 & 15,18,21 & 18 \\
 \hdashline 
\end{array}"



d.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 9 & 1/20 & 9/20 & 81/20 \\
 \hdashline
 10 & 1/20&10/20 & 100/20 \\
 \hdashline
 11 & 2/20 & 22/20 & 242/20 \\
 \hdashline
 12 & 3/20& 36/20 & 432/20 \\
 \hdashline
 13 & 3/20& 39/20 & 507/20 \\
 \hdashline
 14 & 3/20& 42/20 & 588/20 \\
 \hdashline
 15 & 3/20& 45/20 & 675/20 \\
 \hdashline
 16 & 2/20& 32/20 & 512/20 \\
 \hdashline
 17 & 1/20& 17/20 & 289/20 \\
 \hdashline
 18 & 1/20& 18/20 & 324/20 \\
 \hdashline
\end{array}"



Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=\dfrac{270}{20}=13.5=\mu"



The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{3750}{20}-(13.5)^2=5.25= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

"\sigma_{\bar{X}}=\sqrt{5.25}\approx2.291288"